OpenStudy (steve816):

Okay, I need some explanation. For the law of sines, how do you figure out if there is no triangle, 1 triangle or 2 triangles. For example, given that a = 3, b = 2, A = 50°, how many triangles does this form?

imqwerty (imqwerty):

:) 1st we gotta know the cosine formula you know it?

OpenStudy (anonymous):

hello alivejeremy

OpenStudy (alivejeremy):

hello

OpenStudy (steve816):

I haven't learned the cosine formula yet, nor do I think you need to know it to figure this out, but my teacher said to subtract the angle that you find from 180 or something like that. Sorry ahhhh

OpenStudy (alivejeremy):

this might help http://home.windstream.net/okrebs/page93.html

OpenStudy (steve816):

lol nice job google searching @alivejeremy

OpenStudy (steve816):

I have already tried that.

OpenStudy (alivejeremy):

oh

OpenStudy (alivejeremy):

i got one more

OpenStudy (alivejeremy):

http://www.themathpage.com/atrig/law-of-sines.htm

OpenStudy (alivejeremy):

http://2000clicks.com/mathhelp/GeometryLawOfSines.aspx

OpenStudy (steve816):

1/10 rating for you @alivejeremy

OpenStudy (alivejeremy):

what

OpenStudy (steve816):

I don't understand the ambiguous case :( Please...

OpenStudy (hwyl):

LAW OF SINES \(\sf \dfrac{sin~A}{a}=\dfrac{sin~B}{b}=\dfrac{sin~C}{c}\) 

OpenStudy (hwyl):



OpenStudy (hwyl):

Remember that the angles are in degrees, so if you are using a calculator, make sure you set it accordingly. \(\sf \dfrac{sin(30)}{a} = \dfrac{sin(55)}{6} \) either take the reciprocal of both sides which will give \(\sf \dfrac{a}{sin(30)} = \dfrac{6}{sin(55)} \) then solve for a \(\sf \dfrac{a}{sin(30)} \times sin(30) = \dfrac{6}{sin(55)} \times sin(30) \)

OpenStudy (hwyl):

or you can automatically cross-multiply then solve for a

OpenStudy (hwyl):

this is an example, review it and then try your problem

OpenStudy (steve816):

Thanks for your effort @hwyl I will review it :)

OpenStudy (hwyl):

if you run into trouble applying or understanding, let us know or let me know and then we can perhaps break it down further.

OpenStudy (hwyl):

It is important to ask questions and with this example, I am expecting that you would ask a question. I won't point out now since you might figure it out yourself.